OpenStudy (anonymous):

this had been kicking my butt for 20 mins now, a circle has the equation 50x^2+50y^2-20x+25y-23=0 find the center radius and then sketch the graph

OpenStudy (anonymous):

x^2 + y^2 -2x/5 + y/2 - 23/50 = 0 Complete squares?

OpenStudy (anonymous):

did you start off by dividing the whole equation by 50?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

then from there how would i complete the squares? using this equation (x^2 - (2x/5)) + (y^2+(1/2y) = 23/50

OpenStudy (anonymous):

x^2 - 2x/5 + 1/25 + y^2 + y/2 + 1/16 = 23/50 + 1/25 + 1/16 (x-(1/5))^2 + (y + (1/4))^2 = whatever that sum of those fractions is. Now, you have it in the form of (x-a)^2 + (y-b)^2 = r^2 form. Center = (a,b).